Find the current through each battery. 18 Ω 9 V- 3Ω 3 V I3=


1
Expert's answer
2022-03-27T18:41:27-0400

Answer

Current is calculating by using ohms law

"I=\frac{V}{R}=\frac{9}{18}=0.5A\\"

And

"I'=\frac{V}{R}=\frac{3}{3}=1A"